# Learning to (Learn at Test Time): RNNs with Expressive Hidden States

### v3

@@ 1 Introduction
We apply two techniques to make TTT layers more efficient on modern GPUs and TPUs. First, similar to the standard practice of taking gradient steps on mini-batches of sequences during regular training for better parallelism, we use mini-batches of tokens during TTT. Second, we develop a dual form for operations inside each TTT mini-batch. The dual form is equivalent in output to the naive implementation, but trains more than 5×5$\times$5 × faster on our TPUs.
@@
Naive RNN $s_{0}=\texttt{vector()}$() $s_{t}=\sigma\left(\theta_{ss}s_{t-1}+\theta_{sx}x_{t}\right)$ $z_{t}=\theta_{zs}s_{t}+\theta_{zx}x_{t}$ $O(1)$
Self-attention $s_{0}=\texttt{list()}$() $s_{t}=s_{t-1}.\texttt{append}(k_{t},v_{t})$ $z_{t}=V_{t}\texttt{softmax}\left(K^{T}_{t}q_{t}\right)$ $O(t)$
Naive TTT $W_{0}=f.\texttt{params()}$() $W_{t}=W_{t-1}-\eta\nabla\ell(W_{t-1};x_{t})$ $z_{t}=f(x_{t};W_{t})$ $O(1)$
@@ 2.1 TTT as updating a hidden state
with learning rate $\eta$222 For now, consider $W_{0}=0$. We will discuss more sophisticated techniques for initializing $W$ in Subsection 2.7. From the compression point of view, every heuristic needs to decide which input to remember or forget. Our $W$ remembers inputs that produce large gradients – intuitively, inputs that make $W$ learn a lot.
@@ 2.2 Training a network with TTT layers
class TTT_Layer(nn.Module):
@@
class Task(nn.Module):
@@
self.theta_K = nn.Param((d1, d2))
self.theta_V = nn.Param((d1, d2))
self.theta_Q = nn.Param((d1, d2))
@@
Figure 5: Naive implementation of a TTT layer with a linear model and online GD in the style of PyTorch. TTT_Layer can be dropped into a larger network like other sequence modeling layers. Training the network will optimize the parameters of Task in TTT_Layer, because both are subclasses of nn.Module. Since Learner is not a subclass of nn.Module, state.model is updated manually in the inner loop for each call of state.train. For simplicity, we sometimes overload model as model.parameters.
@@ 2.3 Learning a self-supervised task for TTT
Since both $W$ and various $\theta$s appear together in Equation 4, we emphasize again their difference in nature. In the inner loop, only $W$ is optimized, therefore written as an argument of $\ell$; the $\theta$s are "hyper-parameters" of this loss function. In the outer loop, $\theta_{K},\theta_{V},\theta_{Q}$ are optimized alongside $\theta_{\text{rest}}$, and $W$ is merely a hidden state, not a parameter. Figure 5 illustrates this difference with code, where $\theta_{K}$ and $\theta_{V}$ are implemented as parameters of the TTT layer, analogous to the Key and Value parameters of self-attention.
@@ 2.4 Parallelization with mini-batch TTT
Our proposed solution – mini-batch gradient descent – is shown in Figure 6. Denote the TTT batch size by $b$. We use $G_{t}=\nabla\ell\left(W_{t^{\prime}};x_{t}\right)$, where $t^{\prime}=t\,-\,$mod$(t,b)$ is the last timestep of the previous mini-batch (or 0 for the first mini-batch), so we can parallelize $b$ gradient computations at a time. Empirically, $b$ controls a trade-off between speed and quality, as shown in Figure LABEL:fig:batch-size. We chose $b=16$ for all experiments in this paper.
@@
Figure 7: Ablations on TTT mini-batch size $b$, where $b=1$ is online GD and $b=T$ is batch GD. We choose $b=16$ for all experiments in this paper. Left: Smaller $b$ improves perplexity since more GD steps are taken.555 In theory, $b$ can potentially be too small such that the variance between mini-batches is too high, hurting optimization. However, we have not observed such an effect in practice.   The perplexity of 11.09 at $b=16$ corresponds to the final result of TTT-Linear in Figure 10. Right: Forward time in dual form, with context length $T=2048$. Total time (orange) can be decomposed into time for computing the $W$s at the end of every mini-batch (blue) and time for $z_{1},\dots,z_{T}$ (orange -- blue).666 For Figure LABEL:fig:batch-size, we use a single TTT layer in TTT-Linear 1.3B, implemented in pure PyTorch. Our fused kernel significantly improves time efficiency, but makes it difficult to cleanly decompose the time for computing $W_{b}$ vs. $z_{1},\dots,z_{b}$.   Time complexity for the $W$s is $O(T\times d^{2})$, constant in $b$, but the blue line decreases as larger $b$ allows more parallelization until hardware utilization saturates. Time complexity for $z_{1},\dots,z_{T}$ is $O(T\times b\times d)$, so the orange line first decreases with more parallelization, then increases as the extra computation for $z_{1},\dots,z_{T}$ becomes dominant.
@@ 2.5 Dual form
Time complexity of the primal form within a TTT mini-batch is $O(b\times d^{2})$. Time complexity of the dual form is $O(b\times d^{2})$ for computing $W_{b}$ alone, then an additional $O(b^{2}\times d)$ for computing $z_{1},\dots,z_{b}$. Compared to the primal, the dual form sacrifices theoretical complexity for hardware utilization. In practice, $d$ is typically a few hundred and $b$ is chosen to be only 16. As a consequence, wall-clock time for computing $z_{1},\dots,z_{b}$ is relatively small, as observed in the right panel of Figure LABEL:fig:batch-size. In our JAX implementation, training with the dual form is more than 5×5$\times$5 × faster than with primal.
@@ 2.6 Theoretical equivalences
$z_{t}=f\left(\theta_{Q}x_{t};W_{t}\right)=\sum_{s=1}^{t}(\theta_{V}x_{s})(\theta_{K}x_{s})^{T}(\theta_{Q}x_{t}),\vspace{-1ex}$
@@
Linear attn. improved 15.23 --0.68
@@
+ LN and residual in $f$ 14.05 --1.22
+ mini-batch TTT 12.35 --1.70
+ learnable $\eta$ 11.99 --0.36
+ Mamba backbone 11.09 --0.90
@@ 2.7 Implementation details
We propose two variants of TTT layers – TTT-Linear and TTT-MLP, differing only in their instantiations of $f$. For TTT-Linear, $f_{\,\texttt{lin}}(x)=Wx$, where $W$ is square. For TTT-MLP, $f_{\,\texttt{MLP}}$ has two layers similar to the MLPs in Transformers. Specifically, the hidden dimension is 4×4$\times$4 × the input dimension, followed by a GELU activation [29]. For better stability during TTT, $f$ always contains a Layer Normalization (LN) and residual connection. That is, $f(x)=x+\texttt{LN}(f_{\,\texttt{res}}(x)),$ where $f_{\,\texttt{res}}$ can be $f_{\,\texttt{lin}}$ or $f_{\,\texttt{MLP}}$.
@@
The TTT initialization $W_{0}$ is shared between all sequences, even though subsequent weights $W_{1},\dots,W_{T}$ are different for each input sequence. Instead of setting $W_{0}=0$, we can learn it as part of the outer loop. Since outer-loop parameters are always denoted by $\theta$s instead of $W$s, we assign an alias $\theta_{\text{init}}=W_{0}$. In practice, $\theta_{\text{init}}$ adds a negligible amount of parameters comparing to the reconstruction views $\theta_{K},\theta_{Q},\theta_{V}$, because both its input and output are low dimensional. Empirically, we observe that learning $W_{0}$ significantly improves training stability.
@@ 3 Experiments
Following the Mamba paper [25], we perform standard experiments with 2k and 8k context lengths on the Pile [23], a popular dataset of documents for training open-source LLMs [8]. However, the Pile contains few sequences of length greater than 8k [19]. To evaluate capabilities in long context, we also experiment with context lengths ranging from 1k to 32k in 2×2$\times$2 × increments, on a subset of the Pile called Books3, which has been widely used to train LLMs in long context [45, 3].
@@ 3.2 Long context: Books
To evaluate capabilities in long context, we experiment with context lengths ranging from 1k to 32k in 2×2$\times$2 × increments, using a popular subset of the Pile called Books3. The training recipe here is the same as that for Pile.121212 Following the Mamba paper, we always use 0.5M tokens per training batch regardless of the context length. In other words, for context length $T$ we have $0.5\text{M}\,/\,T$ sequences per batch (assume divisible). From the subset of results in Figure 11, we make a few observations:
@@ 4.1.1 Test-Time Training
4.1.2 Fast Weights
The general idea of fast weights is to update the parameters of a “fast” model on only the most relevant data, as opposed to the conventional practice of updating a “slow” model on all data [67]. This idea has existed since the 1980s [30]. The most relevant data can be the test instance itself, therefore TTT can be viewed as a special case of fast weights. Compared to fast weights, TTT embraces the idea of formulating an explicit learning problem, where the test instance is the target of generalization. Our update rule is also an explicit step of optimization.
The idea of fast weight programmers (FWPs) is to update the fast weights with a “slow” model [58]. As a modern example for language modeling, Clark et al. [15] give a Transformer a final layer of fast weights, whose initialization is trained as slow weights. Our inner-loop weights $W$ can be viewed as “fast” and outer-loop weights $\theta$ as “slow”. Therefore, networks containing TTT layers can be viewed as a special case of FWPs [39], similar to how TTT can be viewed as a special case of fast weights.
Modern RNN layers such as linear attention [37, 56] and DeltaNet [55, 75] are inspired by the idea of FWPs. Given their relevance to our work, we discuss these modern RNN layers in detail in the next subsection.
@@ 5 Discussion
  • Multi-level learning to learn. If $f$ itself is a self-attention layer, then by Theorem 2 it can be interpreted as yet another inner loop nested inside the existing one. In this fashion, we can potentially build many levels of nested learning problems.
@@ Appendix B Nadaraya-Watson estimator
Table 3: Training configurations for all experiments. This table reproduces Table 12 in the Mamba paper. The only difference is that the learning rate they use for Mamba and Transformer is 5×5$\times$5 × the values in their Table 12, and we report the actual values (5×5$\times$5 ×). Note that this table only applies to TTT-Linear, TTT-MLP, and Transformers, as Mamba does not follow the multi-head residual block structure inherited from Transformers.
@@ Appendix C Experiment details
  • Cosine schedule: decay to end learning rate $51e-$ 5
@@
To ensure fairness to Mamba, all methods in these experiments have matched training FLOPs and are trained with the same recipe (last row of Table 3) as Mamba 1.4B. For TTT-Linear and TTT-MLP, matched training FLOPs also imply matched inference FLOPs. Transformer (TF finetune) has 2.8×2.8$\times$2.8 × the inference FLOPs, giving it an advantage as our baseline. To match training FLOPs with Mamba, Transformer has 19 blocks instead of 24. For TTT-Linear and TTT-MLP, their training FLOPs are already close to those of Mamba, so we only need to change the hidden dimension of the MLP blocks from 5504 to 5808 for TTT-Linear and 5248 for TTT-MLP.
@@
Figure 14: The self-supervised TTT loss $\ell$ averaged over all test sequences of the form $x_{1},\dots,x_{T}$ where $T=2048$, for all 12 TTT layers in a network with 125M parameters train on the Pile. The same network is also used for $b=1$ (online GD) in the left panel of Figure LABEL:fig:batch-size. For layers in the middle, we observe that $\|x_{t}\|$ rises steadily, causing all three losses to rise with it. Even for these layers, the gap between $\ell(W_{0};x_{t})$ and $\ell(W_{t};x_{t})$ still increases with $t$ For visual clarity, loss values have been averaged over a sliding window of 10 timesteps.

### v4

@@ 1 Introduction
We apply two techniques to make TTT layers more efficient on modern GPUs and TPUs. First, similar to the standard practice of taking gradient steps on mini-batches of sequences during regular training for better parallelism, we use mini-batches of tokens during TTT. Second, we develop a dual form for operations inside each TTT mini-batch. The dual form is equivalent in output to the naive implementation, but trains more than $5\times$ faster on our TPUs.
@@
Naive RNN $s_{0}=\texttt{vector()}$ $s_{t}=\sigma\left(\theta_{ss}s_{t-1}+\theta_{sx}x_{t}\right)$ $z_{t}=\theta_{zs}s_{t}+\theta_{zx}x_{t}$ $O(1)$
Self-attention $s_{0}=\texttt{list()}$ $s_{t}=s_{t-1}.\texttt{append}(k_{t},v_{t})$ $z_{t}=V_{t}\texttt{softmax}\left(K^{T}_{t}q_{t}\right)$ $O(t)$
Naive TTT $W_{0}=f.\texttt{params()}$ $W_{t}=W_{t-1}-\eta\nabla\ell(W_{t-1};x_{t})$ $z_{t}=f(x_{t};W_{t})$ $O(1)$
@@ 2.1 TTT as updating a hidden state
with learning rate $\eta$.222 For now, consider $W_{0}=0$. We will discuss more sophisticated techniques for initializing $W$ in Subsection 2.7. From the compression point of view, every heuristic needs to decide which input to remember or forget. Our $W$ remembers inputs that produce large gradients – intuitively, inputs that make $W$ learn a lot.
@@ 2.2 Training a network with TTT layers
class TTT_Layer($n$.Module):
@@
class Task($n$.Module):
@@
self.theta_K = $n$.Param((d1, d2))
self.theta_V = $n$.Param((d1, d2))
self.theta_Q = $n$.Param((d1, d2))
@@
Figure 5: Naive implementation of a TTT layer with a linear model and online GD in the style of PyTorch. TTT_Layer can be dropped into a larger network like other sequence modeling layers. Training the network will optimize the parameters of Task in TTT_Layer, because both are subclasses of $n$.Module. Since Learner is not a subclass of $n$.Module, state.model is updated manually in the inner loop for each call of state.train. For simplicity, we sometimes overload model as model.parameters.
@@ 2.3 Learning a self-supervised task for TTT
[…]
@@ 2.4 Parallelization with mini-batch TTT
Our proposed solution – mini-batch gradient descent – is shown in Figure 6. Denote the TTT batch size by $b$. We use $G_{t}=\nabla\ell\left(W_{t^{\prime}};x_{t}\right)$, where $t^{\prime}=t\,$-\,mod(t,b)(t,b) is the last timestep of the previous mini-batch (or 0 for the first mini-batch), so we can parallelize $b$ gradient computations at a time. Empirically, $b$ controls a trade-off between speed and quality, as shown in Figure LABEL:fig:batch-size. We chose $b=16$ for all experiments in this paper.
@@
Figure 7: Ablations on TTT mini-batch size $b$, where $b=1$ is online GD and $b=T$ is batch GD. We choose $b=16$ for all experiments in this paper. Left: Smaller $b$ improves perplexity since more GD steps are taken.555 In theory, $b$ can potentially be too small such that the variance between mini-batches is too high, hurting optimization. However, we have not observed such an effect in practice.   The perplexity of 11.09 at $b=16$ corresponds to the final result of TTT-Linear in Figure 10. Right: Forward time in dual form, with context length $T=2048$. Total time (orange) can be decomposed into time for computing the $W$s at the end of every mini-batch (blue) and time for $z_{1},\dots,z_{T}$ (orange - blue).666 For Figure LABEL:fig:batch-size, we use a single TTT layer in TTT-Linear 1.3B, implemented in pure PyTorch. Our fused kernel significantly improves time efficiency, but makes it difficult to cleanly decompose the time for computing $W_{b}$ vs. $z_{1},\dots,z_{b}$.   Time complexity for the $W$s is $O(T\times d^{2})$, constant in $b$, but the blue line decreases as larger $b$ allows more parallelization until hardware utilization saturates. Time complexity for $z_{1},\dots,z_{T}$ is $O(T\times b\times d)$, so the orange line first decreases with more parallelization, then increases as the extra computation for $z_{1},\dots,z_{T}$ becomes dominant.
@@ 2.5 Dual form
Time complexity of the primal form within a TTT mini-batch is $O(b\times d^{2})$. Time complexity of the dual form is $O(b\times d^{2})$ for computing $W_{b}$ alone, then an additional $O(b^{2}\times d)$ for computing $z_{1},\dots,z_{b}$. Compared to the primal, the dual form sacrifices theoretical complexity for hardware utilization. In practice, $d$ is typically a few hundred and $b$ is chosen to be only 16. As a consequence, wall-clock time for computing $z_{1},\dots,z_{b}$ is relatively small, as observed in the right panel of Figure LABEL:fig:batch-size. In our JAX implementation, training with the dual form is more than $5\times$ faster than with primal.
@@ 2.6 Theoretical equivalences
$z_{t}=f\left(\theta_{Q}x_{t};W_{t}\right)=\sum_{s=1}^{t}(\theta_{V}x_{s})(\theta_{K}x_{s})^{T}(\theta_{Q}x_{t}),\vskip-4.30554pt$
@@
Linear attn. improved 15.23 -0.68
@@
+ LN and residual in $f$ 14.05 -1.22
+ mini-batch TTT 12.35 -1.70
+ learnable $\eta$ 11.99 -0.36
+ Mamba backbone 11.09 -0.90
@@ 2.7 Implementation details
We propose two variants of TTT layers – TTT-Linear and TTT-MLP, differing only in their instantiations of $f$. For TTT-Linear, $f_{\,\texttt{lin}}(x)=Wx$, where $W$ is square. For TTT-MLP, $f_{\,\texttt{MLP}}$ has two layers similar to the MLPs in Transformers. Specifically, the hidden dimension is $4\times$ the input dimension, followed by a GELU activation [31]. For better stability during TTT, $f$ always contains a Layer Normalization (LN) and residual connection. That is, $f(x)=x+\texttt{LN}(f_{\,\texttt{res}}(x)),$ where $f_{\,\texttt{res}}$ can be $f_{\,\texttt{lin}}$ or $f_{\,\texttt{MLP}}$.
@@
[…]
@@ 3 Experiments
Following the Mamba paper [27], we perform standard experiments with 2k and 8k context lengths on the Pile [24], a popular dataset of documents for training open-source LLMs [8]. However, the Pile contains few sequences of length greater than 8k [19]. To evaluate capabilities in long context, we also experiment with context lengths ranging from 1k to 32k in $2\times$ increments, on a subset of the Pile called Books3, which has been widely used to train LLMs in long context [52, 3].
@@ 3.2 Long context: Books
To evaluate capabilities in long context, we experiment with context lengths ranging from 1k to 32k in $2\times$ increments, using a popular subset of the Pile called Books3. The training recipe here is the same as that for Pile.121212 Following the Mamba paper, we always use 0.5M tokens per training batch regardless of the context length. In other words, for context length $T$ we have $0.5\text{M}\,/\,T$ sequences per batch (assume divisible). From the subset of results in Figure 11, we make a few observations:
@@ 4.1.1 Test-Time Training
4.1.2 Fast Weights and Fast Weight Programmers
The general idea of fast weights is to update the parameters of a “fast” model on only the most relevant data, as opposed to the conventional practice of updating a “slow” model on all data [74]. This idea has existed since the 1980s [20, 32, 78]. The most relevant data often includes the test instance itself, therefore TTT can be viewed as a special case of fast weights. Compared to fast weights, TTT embraces the idea of formulating an explicit learning problem, where the test instance is the target of generalization. Our update rule is also an explicit step of optimization.
The idea of fast weight programmers (FWPs) is to update the fast weights at test time with a “slow” model that is updated less frequently, if at all [65]. Our inner-loop weights $W$ can be viewed as “fast” and outer-loop weights $\theta$ as “slow”. Therefore, networks containing TTT layers can be viewed as a special case of FWPs [46], similar to how TTT can be viewed as a special case of fast weights. Notably, one instantiation in Irie et al. [38] makes the fast weights an MLP, preceding our TTT-MLP.
Many other modern RNN layers such as linear attention [44, 63] and DeltaNet [62, 83] are also inspired by the idea of FWPs. Given their relevance to our work, we discuss these modern RNN layers in detail in the next subsection. For the rest of this subsection, we briefly outline a few other forms of FWPs. Clark et al. [15] give a Transformer a final layer of fast weights, whose initialization is trained as slow weights. Irie et al. [39] design the fast weights to be programmed by themselves, which can be interpreted as recursive self-improvement. In addition, [40] builds an image generator using the images as fast weights, [37] applies continuous-time extensions of FWPs to time-series classification, while [36] and [26] demonstrate how the choice of update rules affects the expressiveness of FWPs on formal language recognition tasks.
@@ 5 Discussion
  • Multi-level learning to learn. If $f$ itself is a self-attention layer, then by Theorem 2 it can be interpreted as yet another inner loop nested inside the existing one. In this fashion, we can potentially build many levels of nested learning problems. Versions of this idea have already been explored in [38] and [39].
@@ Appendix B Nadaraya-Watson estimator
Table 3: Training configurations for all experiments. This table reproduces Table 12 in the Mamba paper. The only difference is that the learning rate they use for Mamba and Transformer is $5\times$ the values in their Table 12, and we report the actual values ($5\times$). Note that this table only applies to TTT-Linear, TTT-MLP, and Transformers, as Mamba does not follow the multi-head residual block structure inherited from Transformers.
@@ Appendix C Experiment details
  • Cosine schedule: decay to end learning rate 1e51e-5
@@
To ensure fairness to Mamba, all methods in these experiments have matched training FLOPs and are trained with the same recipe (last row of Table 3) as Mamba 1.4B. For TTT-Linear and TTT-MLP, matched training FLOPs also imply matched inference FLOPs. Transformer (TF finetune) has $2.8\times$ the inference FLOPs, giving it an advantage as our baseline. To match training FLOPs with Mamba, Transformer has 19 blocks instead of 24. For TTT-Linear and TTT-MLP, their training FLOPs are already close to those of Mamba, so we only need to change the hidden dimension of the MLP blocks from 5504 to 5808 for TTT-Linear and 5248 for TTT-MLP.
@@
Figure 14: The self-supervised TTT loss $\ell$ averaged over all test sequences of the form $x_{1},\dots,x_{T}$ where $T=2048$, for all 12 TTT layers in a network with 125M parameters train on the Pile. The same network is also used for $b=1$ (online GD) in the left panel of Figure LABEL:fig:batch-size. For layers in the middle, we observe that $\|x_{t}\|$ rises steadily, causing all three losses to rise with it. Even for these layers, the gap between $\ell(W_{0};x_{t})$ and $\ell(W_{t};x_{t})$ still increases with $t$ . For visual clarity, loss values have been averaged over a sliding window of 10 timesteps.